OpenStudy (anonymous):

17,19,20,22,24,24,28,30,30,36 what is the mean?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

because 24 and 24 are right next to eachother you take 24+24/2=24

myininaya (myininaya):

all you do is just add the numbers together and divide that sum by the number of numbers there are

OpenStudy (anonymous):

thats mean i believe this is a stats question

myininaya (myininaya):

\[\frac{17+19+20+22+24+24+28+30+30+36}{10}\]

myininaya (myininaya):

it says mean

OpenStudy (anonymous):

haha i thought median ive been blowin it big time myani tonight, big time

myininaya (myininaya):

lol yes that would be the median

myininaya (myininaya):

(24+24)/2=median

OpenStudy (anonymous):

so tell me the answer

OpenStudy (anonymous):

myini had the correct answer just add up all the numbers and divide by 10